Don is twice as old as Irene; Ed is 16 years older than irene. the sum of all three ages is 6-. find the ages of all of them.
Answer by ankor@dixie-net.com(22740)

Don is twice as old as Irene; Ed is 16 years older than Irene.
the sum of all three ages is 60. find the ages of all of them.
:
Let I = Irene's present age
then
2I = Don's present age
and
(I+16) = Ed's age
:
I + 2I + (I+16) = 60
4I = 60 - 16
4I = 44
I = 11 is Irene's present age
then
2(11) = 22 yrs is Don's age
and
11 + 16 = 27 yrs is Ed's age
:
:
11 + 22 + 27 = 60